There are 1200 pupils 45% are boys,how many girls are there?
Found 2 solutions by rm29924, MathTherapy:
Answer by rm29924(97)

45/100=x/1200
we are solving to see what 45 percent of 1200 is
make a ratio 45 out of 100 equals what number out of 1200
then you cross multiply
100x=45(1200)
100x=54000
divide both sides by 100
x=540 that is the number of boys there are
take the total 1200 and subtract the 540 boys
1200-540=660
there are 660 girls

Answer by MathTherapy(10553)


There are 1200 pupils 45% are boys,how many girls are there?

Since 45% of the 1,200 students are boys, then 55% of the 1,200 students are girls.

You should be able to do the calculation!!

Send comments and “thank-yous” to “D” at MathMadEzy@aol.com